The average expenditure per student (based on average daily attendance) for a certain school year was $10,337 with a population standard deviation of $1560. A survey for the next school year of 150 randomly selected students resulted in a sample mean of $10, 798. Find the 95% confidence level? Should the null hypothesis be rejected?



1
Expert's answer
2022-03-04T04:35:57-0500

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96." The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(10798-1.96\times\dfrac{1560}{\sqrt{150}}, 10798+1.96\times\dfrac{1560}{\sqrt{150}})"

"=(10548.35, 11047.65)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "10548.35 < \mu < 11047.65," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(10548.35, 11047.65)."

The following null and alternative hypotheses need to be tested:

"H_0:\mu=10337"

"H_1:\mu\not=10337"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{X}-\mu}{\sigma/\sqrt{n}}=\dfrac{10798-10337}{1560/\sqrt{150}}\approx3.6193"

Since it is observed that "|z| = 3.619 > 1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(Z>3.6193)=0.000295," and since "p=0.000295<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than "10337," at the "\alpha = 0.05" significance level.